Since the LHC restarted after the last technical stop it has added about 210/pb to the integrated luminosity. In the last 5 days since returning to 1380 bunch fills more than 150/pb have been added which brings them back to the previous good production rate of about 30/pb per day.
New records for peak luminosity were set last night with 1.75/nb/s during stable beams and 1.8/ns/s during the adjust mode of a run that was dumped before stable beams. Before the technical stop the record was 1.28/nb/s so significant luminosity increases have already been achieved. This has been mainly due to better beam emittance at about 2.4 um in the lastest runs. The bunch intensity is still being held at 120 billion.
There are many problems causing beams to be dumped including UFOs, cryo, vacuum, RF and power glitches. However it has been possible to recycle and return to stable beams relatively quickly to keep the data flowing. Unless they can improve the stability they may have to reconsider the plan to increase bunch intensity because that would mean more cryogenic outages that can stop the runs for one or two days at a time.
Steve Myers will give a status report at EPS in a plenary session on Wednesday that is due to be webcast.